# Towards Unbiased and Robust Spatio-Temporal Scene Graph Generation and Anticipation

### v1

@@ 4.4 Robustness Evaluation
Fog STTran+ [35] 9.8 10.6 10.6
   +ImparTail (Ours) 11.4 (+16.3%) 12.2 (+15.1%) 12.2 (+15.1%)
DSGDetr+ [35] 10.8 12.0 12.0
   +ImparTail (Ours) 15.2 (+40.7%) 16.6 (+38.3%) 16.6 (+38.3%)
STTran++ [35] 11.7 12.9 12.9
   +ImparTail (Ours) 15.8 (+35.0%) 17.4 (+34.9%) 17.4 (+34.9%)
DSGDetr++ [35] 12.7 13.8 13.8
   +ImparTail (Ours) 15.8 (+24.4%) 17.4 (+26.1%) 17.4 (+26.1%)
@@
5 Experiments
We apply ImparTail to generate and anticipate STSGs on the Action Genome dataset [17]. We prepare the dataset by following the pre-processing step described in [7, 35]. The dataset includes 35 object classes and 25 relationship classes. These 25 relationship classes are grouped into three categories, namely: (1) Attention Relations comprise relationship classes which primarily describe the attention of the subject towards the object, (2) Spatial Relations comprise relationship classes that describe the spatial relationship between the subject and objects, and (3) Contacting Relations comprises relationship classes that indicate different ways the subject contacts the object  [^6] Remark. The dataset predominantly features videos with a single individual interacting with objects across diverse real-world settings. Within the Action Genome framework, a subject-object pair can exhibit multiple types of spatial and contact relationships..
In the following sections, we will assess the performance of ImparTail on four tasks, namely: (a) VidSGG, (b) SGA, (c) RobustVidSGG and (d) RobustSGA. For each task, we first select a few methods, train and evaluate them as per the instructions, then adapt them to ImparTail framework, re-train and evaluate them. For VidSGG and Robust VidSGG tasks, we selected STTran [7] and DSGDetr [10] as our baselines and for the SGA task, we selected STTran+ [35], STTran++ [35], DSGDetr+ [35], DSGDetr++ [35], SceneSayerODE [35], SceneSayerSDE [35] as our baselines.
Settings.
We follow [7] and use three different settings to evaluate models for the VidSGG task. Each setting has a different level of information provided as input to the models. (a) Scene Graph Detection (SGDET): We only provide frames of the video as input, (b) Scene Graph Classification (SGCLS): Along with frames, bounding box information is provided as input to the model and (c) Predicate Classification (PREDCLS): We provide frames, bounding boxes and object labels as input to the model. Similarly, for the SGA task, we follow SceneSayer [35] and use three different settings to evaluate models. (a) Action Genome Scenes (AGS): In AGS, the model’s input is limited to raw frames of the video. (b) Partially Grounded Action Genome Scenes (PGAGS): In this intermediate setting, along with raw video frames, we additionally input the model with precise bounding box information of active interacting objects observed in the scene. (c) Grounded Action Genome Scenes (GAGS): In this model takes precise bounding box information and the categories of the observed interacting objects as input.
5.1 Video Scene Graph Generation.
We began by training and evaluating the selected baseline models. Subsequently, we adapted these models to incorporate ImparTail, followed by retraining and evaluation. Our experiments encompassed three VidSGG modes SGDET, SGCLS, and PREDCLS and were reported under three distinct settings: (1) With Constraint, (2) No Constraint, and (3) Semi Constraint. As demonstrated in Tab 1, models trained using ImparTail consistently outperformed baselines.
[…]
@@ 5 Experiments
To train and evaluate the SGA models, we adopted a methodology similar to that of VidSGG. Initially, we trained and assessed the selected baseline models. Subsequently, we integrated these models into the ImparTail framework, followed by retraining and re-evaluating their performance. As demonstrated in Table 2, Models trained using ImparTail consistently outperformed their baseline variants.
@@
We evaluated VidSGG and SGA models trained with ImparTail and those trained using traditional approaches across 16 input corruption scenarios. Our experiments revealed that introducing noise to the inputs generally led to a decline in performance. However, models trained with ImparTail exhibited a less severe performance drop than baseline models. For example, in the PREDCLS mode under the No Constraint setting with Gaussian Noise for STTran, the baseline model’s mR@10 decreased from 45.7 to 31.4. In contrast, the ImparTail-trained model only dropped from 65.5 to 57.5. To maintain brevity, only a subset of the results are reported in the main paper, and we defer additional results to the supplementary text.
Insights: (1) We noticed that the drop in performance of both VidSGG and SGA models trained using ImparTail is lower than the models trained with traditional approaches. (2) as illustrated in Tables 3, 4, our models performed significantly better than the baseline variants, (3) We also observed that models trained using ImparTail not only exhibited consistent strong mean recall scores but also outperformed others in recall metrics sometimes (refer Tab 3).

### v2

@@ 5 Experiments
Settings. We follow [7] and use three different settings to evaluate models for the VidSGG task. Each setting has a different level of information provided as input to the models. (a) Scene Graph Detection (SGDET): We only provide frames of the video as input, (b) Scene Graph Classification (SGCLS): Along with frames, bounding box information is provided as input to the model and (c) Predicate Classification (PREDCLS): We provide frames, bounding boxes and object labels as input to the model. Similarly, for the SGA task, we follow SceneSayer [37] and use three different settings to evaluate models. (a) Action Genome Scenes (AGS): In AGS, the model’s input is limited to raw frames of the video. (b) Partially Grounded Action Genome Scenes (PGAGS): In this intermediate setting, along with raw video frames, we additionally input the model with precise bounding box information of active interacting objects observed in the scene. (c) Grounded Action Genome Scenes (GAGS): In this model takes precise bounding box information and the categories of the observed interacting objects as input.
Remark. Although ImparTail is designed to learn estimators ranging from biased to unbiased through tuning the mask ratio in mask generation algorithm Alg 3, our experiments on ActionGenome revealed that a fixed Stochastic Masking Ratio of 0.9 delivered similar performance in mean recall metrics to a sequential strategy that started at a masking ratio of 0.6 and gradually increased to 0.9. Moreover, the fixed ratio also showed competitive results within fewer training epochs. Therefore, in the paper to report results, we opted for a fixed Masking Ratio of 0.9999Due to the skewed distribution of the dataset, using a 0.9 masking ratio creates a scenario where, in each training epoch, the model primarily samples head (easier) classes while leaving tail (harder) classes unaltered.. For completeness, we provide a thorough ablation in the Appendix.
5.1 Video Scene Graph Generation.
We first trained and evaluated the chosen baseline models. Next, we integrated ImparTailinto these baselines and conducted retraining and evaluation. Our experiments encompassed three VidSGG modes SGDET, SGCLS, and PREDCLS and each included three distinct settings: (1) With Constraint, (2) No Constraint, and (3) Semi Constraint. As demonstrated in Table 1, models trained using ImparTail consistently outperformed the chosen VidSGG baselines.
[…]
@@
We employed a procedure for training and evaluating SGA models that closely mirrors our approach for VidSGG. First, we trained and evaluated the chosen baseline models. We then integrated these models into the ImparTail framework, retrained them, and re-evaluated their performance. As demonstrated in Table 2, Models trained using ImparTail consistently outperformed their baseline variants.
@@
We evaluated VidSGG and SGA models trained with ImparTail and those trained using traditional approaches across 16 input corruption scenarios. Our experiments revealed that introducing noise to the inputs generally led to a decline in performance. However, models trained with ImparTail exhibited a less severe performance drop than baseline models. For example, in the PREDCLS mode under the No Constraint setting with Gaussian Noise for STTran, the baseline model’s mR@10 decreased from 45.7 to 31.4. In contrast, the ImparTail-trained model dropped from 65.5 to 57.5.
Figure 5: Predicate Classification recall performance. of models trained using existing VidSGG methods STTran, DSGDetr, and models trained using their ImparTail adaptations. In each row, we compare the R@10/50 performance of each relationship category without corruptions(left) and with corruptions(right) in input data).
Insights: (1) We noticed that the drop in performance of both VidSGG and SGA models trained using ImparTail is lower than the models trained with traditional approaches (see Tables 3, 4). (2) We also observed that models trained using ImparTail not only exhibited consistent strong mean recall scores but also outperformed baselines in recall metrics sometimes (refer Tab 3). (3) ImparTail yields models that exhibit: (a) Comprehensive Relational Knowledge: ImparTail captures when certain relationships rarely co-occur (e.g., (sitting_on, in_front_of)) and when they often appear together (e.g., (sitting_on, beneath), (sitting_on, behind)). Thus, ImparTail trained models focus on a broader spatio-temporal context and other discovered associations to infer missing connections when frames are corrupted. (b) Better Generalization: By balancing out the data distribution, ImparTail avoids overfitting to the most frequent object-relationship patterns (refer to Figure 5). This regularizing effect (similar in spirit to dropout) pushes the model to rely on shared patterns across categories, not just the apparent cues in dominant categories. Consequently, when inputs are noisy, the model is more likely to recover the underlying spatio-temporal relationships using generalized features.